OpenStudy (anonymous):

Which rectangular equation corresponds to these parametric equations? x = 0.2sec t and y = -0.25tan t

OpenStudy (perl):

what do we know about tan and sec, there is a pythagorean identity

OpenStudy (anonymous):

do you guys mind helping me when your done here?

OpenStudy (anonymous):

eliminate t using trig identities. 1+tan^2 t = sec^2 t. tan t=-4y, sec t = 5x, so 1+16y^2=25x^2

OpenStudy (perl):

sin^2(u) + cos^2(u) = 1 tan^2(u) + 1 = sec^2(u) 1 + cot^2(u) = csc^2(u)

OpenStudy (anonymous):

soooo.... your answer is 25x^2 − 16y^2 = 1

OpenStudy (anonymous):

yep

OpenStudy (freckles):

well I think @Niguyver it is the one where you have \[1=\sec^2(t)-\tan^2(t)\] (this is the Pythagorean identity that was talked about above) solve for sec(t) and tan(t) given your equations then square and plug in

OpenStudy (freckles):

should be no + between the sec^2 and tan^2